In economy, the consumer expenditure is represented by the following equation:

C = 50 + 0.75 Yd

Assume that investors spend RM500 at every level of income ( I = RM500), net exports are zero
(X – M) = 0, goverment purchases are RM300 and taxes are RM200.

a) Determine the equilibrium level of GDP.
(4)

b) If potential GDP is RM3,000, is there a recessionary or inflationary gap? If so, how much?
(2)

c) What will happen to the equilibrium level of GDP if investors become optimistic about the country’s future and raise their investment to RM600?
(4)

d) After investment has increased to RM600, is there a recessionary or inflationary gap? How much?
(2)
1
Expert's answer
2015-07-22T00:00:44-0400
C = 50 + 0.75 Yd
I = RM500, (X – M) = 0, G = RM300 and taxes are RM200.
a) The equilibrium level of GDP is:
Y = C + I + G + NX
Y = 50 + 0.75(Y - 200) + 500 + 300 + 0
0.25Y = 700
Y = RM2800
b) If potential GDP is RM3,000, there is a recessionary gap of RM200.
c) If investors become optimistic about the country’s future and raise their investment to RM600, the equilibrium level of GDP will be:
Y = 50 + 0.75(Y - 200) + 600 + 300 + 0
0.25Y = 800
Y = RM3200
d) After investment has increased to RM600 there is an inflationary gap of RM200.